# Commute Graph Neural Networks

### v3

@@ 5.4 CGNN
The randomized truncated SVD to compute $\mathcal{R}^{\dagger}$ is $\mathcal{O}(q|E|)$ where $q$ is the required rank, and the message passing iteration has the same time complexity as DirGNN with $\mathcal{O}(L|E|d^{2})$. Therefore, the overall time complexity of CGNN is $\mathcal{O}((Ld^{2}+q)|E|)$. In practice, $q$ is typically set to 5, rendering the time complexity effectively linear with respect to the number of edges $|E|$. In GNN domain, models with a complexity less than $\mathcal{O}(N^{2})$ are generally considered feasible by researchers [34]. Given that real-world networks are often extremely sparse, i.e., $|E|\ll N^{2}$, CGNN demonstrates its feasibility as a model within the GNN family.
@@ 6.1 Experimental Settings
We conduct node classification experiments on eight digraph datasets. We provide a performance comparison with 12 baselines including 1) General GNNs: GCN [35], GAT [36], and GraphSAGE [21]; 2) Non-local GNNs: APPNP [37], MixHop [38], GPRGNN [39], and GCNII [40]; 3) Digraph NNs: DGCN [7], DiGCN [6], MagNet [9], DiGCL [8], DUPLEX [23], and DirGNN [10]. For all baselines, we apply both the symmetrized and asymmetric adjacency matrix for experiments. The results reported are the better of the two results.
We evaluate the performance by node classification accuracy with standard deviation in the semi-supervised setting. For Squirrel and Chameleon, we use 10 public splits (48%/32%/20% for training/validation/testing) provided by [41]. For the remaining datasets, we adopt the same splits as [6, 8], which chooses 20 nodes per class for the training set, 500 for the validation set, and allocates the rest to the test set. We conduct our experiments on 2 Intel Xeon Gold 5215 CPUs and 1 NVIDIA GeForce RTX 3090 GPU.
@@ 6.2 Data Statistics
The datasets used in Section 6 are Squirrel, Chameleon [42], Citeseer [43], CoraML [44], AM-Photo [45], Snap-Patents, Roman-Empire, and Arxiv-Year [10]. We summarize their statistics in Table 1. $\mathrm{homo\_ratio}$ represents the homophily ratio, a metric proposed by [46]. which is employed to gauge the degree of homophily within the graph. A lower $\mathrm{homo\_ratio}$ signifies a greater degree of heterophily, indicating a higher prevalence of edges that connect nodes of differing classes.
@@
Method Squirrel Chameleon Citeseer CoraML AM-Photo Snap-Patents Roman-Empire Arxiv-Year GCN 52.43$\pm$2.01 67.96$\pm$1.82 66.03$\pm$1.88 70.92$\pm$0.39 88.52$\pm$0.47 51.02$\pm$0.06 73.69$\pm$0.74 46.02$\pm$0.26 GAT 40.72$\pm$1.55 60.69$\pm$1.95 65.58$\pm$1.39 72.22$\pm$0.57 88.36$\pm$1.25 OOM 49.18$\pm$1.35 45.30$\pm$0.23 GraphSAGE 41.61$\pm$0.74 62.01$\pm$1.06 66.81$\pm$1.38 74.16$\pm$1.55 89.71$\pm$0.57 67.45$\pm$0.53 86.37$\pm$0.80 55.43$\pm$0.75 APPNP 51.91$\pm$0.56 45.37$\pm$1.62 66.90$\pm$1.82 70.31$\pm$0.67 87.43$\pm$0.98 51.23$\pm$0.54 72.96$\pm$0.38 50.31$\pm$0.42 MixHop 43.80$\pm$1.48 60.50$\pm$2.53 56.09$\pm$2.08 65.89$\pm$1.50 87.17$\pm$1.34 41.22$\pm$0.19 50.76$\pm$0.14 45.30$\pm$0.26 GPRGNN 50.56$\pm$1.51 66.31$\pm$2.05 61.74$\pm$1.87 73.31$\pm$1.37 90.23$\pm$0.34 40.19$\pm$0.03 64.85$\pm$0.27 45.07$\pm$0.21 GCNII 38.47$\pm$1.58 63.86$\pm$3.04 58.32$\pm$1.93 64.84$\pm$0.71 83.40$\pm$0.79 48.09$\pm$0.09 74.27$\pm$0.13 57.36$\pm$0.17 DGCN 37.16$\pm$1.72 50.77$\pm$3.31 66.37$\pm$1.93 75.02$\pm$0.50 87.74$\pm$1.02 OOM 51.92$\pm$0.43 OOM DiGCN 33.44$\pm$2.07 50.37$\pm$4.31 64.99$\pm$1.72 77.03$\pm$0.70 88.66$\pm$0.51 OOM 52.71$\pm$0.32 48.37$\pm$0.19 MagNet 39.01$\pm$1.93 58.22$\pm$2.87 65.04$\pm$0.47 76.32$\pm$0.10 86.80$\pm$0.65 OOM 88.07$\pm$0.27 60.29$\pm$0.27 DUPLEX 57.60$\pm$0.98 61.25$\pm$0.94 67.60$\pm$0.72 72.26$\pm$0.71 87.80$\pm$0.82 66.54$\pm$0.11 79.02$\pm$0.08 64.37$\pm$0.27 DiGCL 35.82$\pm$1.73 56.45$\pm$2.77 67.42$\pm$0.14 77.53$\pm$0.14 89.41$\pm$0.11 70.65$\pm$0.07 87.94$\pm$0.10 63.10$\pm$0.06 DirGNN 75.19$\pm$1.26 79.11$\pm$2.28 66.57$\pm$0.74 75.33$\pm$0.32 88.09$\pm$0.46 73.95$\pm$0.05 91.23$\pm$0.32 64.08$\pm$0.26 CGNN 77.61$\pm$1.87 79.54$\pm$1.82 70.27$\pm$0.66 77.06$\pm$0.67 90.41$\pm$0.32 72.33$\pm$0.16 92.89$\pm$0.51 66.97$\pm$0.15
@@ 6.3 Overall Results and Analysis
Table 2 reports the node classification results across eight digraph datasets. Our method CGNN achieves new state-of-the-art results on 6 out of 8 datasets, and comparable results on the remaining ones, validating the superiority of CGNN. We provide more observations as follows. Firstly, while non-local GNNs have the potential to cover the commute paths between adjacent nodes by stacking multiple layers, they do not consistently outperform general, shallow GNN models. It suggests that coarsely aggregating all nodes in commute paths is ineffective. The reason is that deeper models may aggregate excessive irrelevant information for the central node. Our goal is to encode mutual relationships between adjacent nodes by considering their commute times. Aggregating all nodes along the entire path introduces excessive information about other nodes unrelated to the direct relationship between the target nodes. Secondly, GNNs tailored for digraphs do not seem to bring substantial gains. Our results show that with careful hyper-parameter tuning, general GNNs can achieve results comparable to, or even better than, those tailored for digraphs, as evidenced in the Squirrel, Chameleon, and AM-Photo datasets. Thirdly, CGNN achieves state-of-the-art results on both homophilic and heterophilic digraph benchmarks. Notably, DirGNN also performs comparably on heterophilic graphs (e.g., Squirrel and Chameleon), supporting the findings of [10] that distinguishing edge directionality during message passing enables the central node to adaptively balance information flows from both heterophilic and homophilic neighbors, effectively mitigating the impact of heterophily. Moreover, CGNN, an enhanced version of DirGNN, further improves performance on these graphs by effectively incorporating commute times to refine the strength of relationships between nodes, enhancing model robustness under heterophily.
@@
Squirrel Chameleon Citeseer CoraML AM-Photo Method Acc. Time Acc. Time Acc. Time Acc. Time Acc. Time CGNN 77.61 102.93 79.54 112.24 70.27 92.26 77.02 127.15 90.41 121.80 CGNNppr 68.37 257.84 71.69 253.05 68.59 137.82 76.23 192.09 88.52 203.04

### v4

@@ 5.4 CGNN
The randomized truncated SVD to compute $\mathcal{R}^{\dagger}$ is $\mathcal{O}(q|E|)$ where $q$ is the required rank, and the message passing iteration has the same time complexity as DirGNN with $\mathcal{O}(L|E|d^{2})$. Therefore, the overall time complexity of CGNN is $\mathcal{O}((Ld^{2}+q)|E|)$. In practice, $q$ is typically set to 5, rendering the time complexity effectively linear with respect to the number of edges $|E|$. In GNN domain, models with a complexity less than $\mathcal{O}(N^{2})$ are generally considered feasible by researchers (Wu et al., 2020). Given that real-world networks are often extremely sparse, i.e., $|E|\ll N^{2}$, CGNN demonstrates its feasibility as a model within the GNN family.
@@ 6.1 Experimental Settings
We conduct node classification experiments on eight digraph datasets. We provide a performance comparison with 12 baselines including 1) General GNNs: GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), and GraphSAGE (Hamilton et al., 2017); 2) Non-local GNNs: APPNP (Klicpera et al., 2019), MixHop (Abu-El-Haija et al., 2019), GPRGNN (Chien et al., 2021), and GCNII (Ming Chen et al., 2020); 3) Digraph NNs: DGCN (Tong et al., 2020b), DiGCN (Tong et al., 2020a), MagNet (Zhang et al., 2021), DiGCL (Tong et al., 2021), DUPLEX (Ke et al., 2024), and DirGNN (Rossi et al., 2023). For all baselines, we apply both the symmetrized and asymmetric adjacency matrix for experiments. The results reported are the better of the two results.
We evaluate the performance by node classification accuracy with standard deviation in the semi-supervised setting. For Squirrel and Chameleon, we use 10 public splits (48%/32%/20% for training/validation/testing) provided by (Pei et al., 2019). For the remaining datasets, we adopt the same splits as (Tong et al., 2020a, 2021), which chooses 20 nodes per class for the training set, 500 for the validation set, and allocates the rest to the test set. We conduct our experiments on 2 Intel Xeon Gold 5215 CPUs and 1 NVIDIA GeForce RTX 3090 GPU.
@@ 6.2 Data Statistics
The datasets used in Section 6 are Squirrel, Chameleon (Rozemberczki et al., 2021), citepseer (Sen et al., 2008), CoraML (Bojchevski & Günnemann, 2017), AM-Photo (Shchur et al., 2018), Snap-Patents, Roman-Empire, and Arxiv-Year (Rossi et al., 2023). We summarize their statistics in Table 1. $\mathrm{homo\_ratio}$ represents the homophily ratio, a metric proposed by (Zhu et al., 2020). which is employed to gauge the degree of homophily within the graph. A lower $\mathrm{homo\_ratio}$ signifies a greater degree of heterophily, indicating a higher prevalence of edges that connect nodes of differing classes.
@@
Method Squirrel Chameleon citepseer CoraML AM-Photo Snap-Patents Roman-Empire Arxiv-Year GCN 52.43$\pm$2.01 67.96$\pm$1.82 66.03$\pm$1.88 70.92$\pm$0.39 88.52$\pm$0.47 51.02$\pm$0.06 73.69$\pm$0.74 46.02$\pm$0.26 GAT 40.72$\pm$1.55 60.69$\pm$1.95 65.58$\pm$1.39 72.22$\pm$0.57 88.36$\pm$1.25 OOM 49.18$\pm$1.35 45.30$\pm$0.23 GraphSAGE 41.61$\pm$0.74 62.01$\pm$1.06 66.81$\pm$1.38 74.16$\pm$1.55 89.71$\pm$0.57 67.45$\pm$0.53 86.37$\pm$0.80 55.43$\pm$0.75 APPNP 51.91$\pm$0.56 45.37$\pm$1.62 66.90$\pm$1.82 70.31$\pm$0.67 87.43$\pm$0.98 51.23$\pm$0.54 72.96$\pm$0.38 50.31$\pm$0.42 MixHop 43.80$\pm$1.48 60.50$\pm$2.53 56.09$\pm$2.08 65.89$\pm$1.50 87.17$\pm$1.34 41.22$\pm$0.19 50.76$\pm$0.14 45.30$\pm$0.26 GPRGNN 50.56$\pm$1.51 66.31$\pm$2.05 61.74$\pm$1.87 73.31$\pm$1.37 90.23$\pm$0.34 40.19$\pm$0.03 64.85$\pm$0.27 45.07$\pm$0.21 GCNII 38.47$\pm$1.58 63.86$\pm$3.04 58.32$\pm$1.93 64.84$\pm$0.71 83.40$\pm$0.79 48.09$\pm$0.09 74.27$\pm$0.13 57.36$\pm$0.17 DGCN 37.16$\pm$1.72 50.77$\pm$3.31 66.37$\pm$1.93 75.02$\pm$0.50 87.74$\pm$1.02 OOM 51.92$\pm$0.43 OOM DiGCN 33.44$\pm$2.07 50.37$\pm$4.31 64.99$\pm$1.72 77.03$\pm$0.70 88.66$\pm$0.51 OOM 52.71$\pm$0.32 48.37$\pm$0.19 MagNet 39.01$\pm$1.93 58.22$\pm$2.87 65.04$\pm$0.47 76.32$\pm$0.10 86.80$\pm$0.65 OOM 88.07$\pm$0.27 60.29$\pm$0.27 DUPLEX 57.60$\pm$0.98 61.25$\pm$0.94 67.60$\pm$0.72 72.26$\pm$0.71 87.80$\pm$0.82 66.54$\pm$0.11 79.02$\pm$0.08 64.37$\pm$0.27 DiGCL 35.82$\pm$1.73 56.45$\pm$2.77 67.42$\pm$0.14 77.53$\pm$0.14 89.41$\pm$0.11 70.65$\pm$0.07 87.94$\pm$0.10 63.10$\pm$0.06 DirGNN 75.19$\pm$1.26 79.11$\pm$2.28 66.57$\pm$0.74 75.33$\pm$0.32 88.09$\pm$0.46 73.95$\pm$0.05 91.23$\pm$0.32 64.08$\pm$0.26 CGNN 77.61$\pm$1.87 79.54$\pm$1.82 70.27$\pm$0.66 77.06$\pm$0.67 90.41$\pm$0.32 72.33$\pm$0.16 92.89$\pm$0.51 66.97$\pm$0.15
@@ 6.3 Overall Results and Analysis
Table 2 reports the node classification results across eight digraph datasets. Our method CGNN achieves new state-of-the-art results on 6 out of 8 datasets, and comparable results on the remaining ones, validating the superiority of CGNN. We provide more observations as follows. Firstly, while non-local GNNs have the potential to cover the commute paths between adjacent nodes by stacking multiple layers, they do not consistently outperform general, shallow GNN models. It suggests that coarsely aggregating all nodes in commute paths is ineffective. The reason is that deeper models may aggregate excessive irrelevant information for the central node. Our goal is to encode mutual relationships between adjacent nodes by considering their commute times. Aggregating all nodes along the entire path introduces excessive information about other nodes unrelated to the direct relationship between the target nodes. Secondly, GNNs tailored for digraphs do not seem to bring substantial gains. Our results show that with careful hyper-parameter tuning, general GNNs can achieve results comparable to, or even better than, those tailored for digraphs, as evidenced in the Squirrel, Chameleon, and AM-Photo datasets. Thirdly, CGNN achieves state-of-the-art results on both homophilic and heterophilic digraph benchmarks. Notably, DirGNN also performs comparably on heterophilic graphs (e.g., Squirrel and Chameleon), supporting the findings of (Rossi et al., 2023) that distinguishing edge directionality during message passing enables the central node to adaptively balance information flows from both heterophilic and homophilic neighbors, effectively mitigating the impact of heterophily. Moreover, CGNN, an enhanced version of DirGNN, further improves performance on these graphs by effectively incorporating commute times to refine the strength of relationships between nodes, enhancing model robustness under heterophily.
@@
Squirrel Chameleon citepseer CoraML AM-Photo Method Acc. Time Acc. Time Acc. Time Acc. Time Acc. Time CGNN 77.61 102.93 79.54 112.24 70.27 92.26 77.02 127.15 90.41 121.80 CGNNppr 68.37 257.84 71.69 253.05 68.59 137.82 76.23 192.09 88.52 203.04
@@ 7 Conclusion
Appendix A Proofs and Derivations
A.1 Derivation of DiLap T
Since the digraph is unweighted, then $\mathbf{P}_{ij}=\frac{1}{d^{\text{out}}_{i}}$ and $\frac{1}{d^{\text{out}}_{i}}\mathbf{A}_{i,:}=\mathbf{P}_{i,:}$. We have:
$(\mathbf{T}s)_{i}$ $=\sum_{v_{j}\in\mathcal{N}^{\text{out}}_{i}}\mathbf{P}_{ij}(s_{i}-s_{j})$ (12)
$=\frac{1}{d^{\text{out}}_{i}}s_{i}-\frac{1}{d^{\text{out}}_{i}}\sum_{v_{j}\in\mathcal{N}^{\text{out}}_{i}}s_{j}$
$=(\mathbf{D}^{-1}s)_{i}-\frac{1}{d^{\text{out}}_{i}}\mathbf{A}_{i,:}s$
$=(\mathbf{D}^{-1}s)_{i}-(\mathbf{P}s)_{i}$
$=((\mathbf{D}^{-1}-\mathbf{P})s)_{i}.$
Thus, DiLap $\mathbf{T}$ can be represented as $\mathbf{D}^{-1}-\mathbf{P}$.
A.2 Proof of Lemma 5.1
Proof.
Since the transition matrix $\widetilde{\mathbf{P}}$ is row-stochastic, it follows that $\widetilde{\mathbf{P}}^{t}\mathbf{J}=\mathbf{J}$. In light of Eq. 2 and considering that $\pi$ is stochastic, we have $\mathbf{Z}\mathbf{J}=\bm{0}_{n\times n}$. Let $\mathcal{K}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\widetilde{\mathcal{T}}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}(\widetilde{\mathbf{D}}^{-1}-\widetilde{\mathbf{P}})\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}$, $\mathcal{J}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}\mathbf{J}\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}$, and $\mathcal{Z}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}\mathbf{Z}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}$, incorporating these into Eq. 3, we have:
$\mathcal{Z}+\mathcal{J}=(\mathcal{K}+\mathcal{J}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})^{-1}.$ (13)
By post-multiplying Eq. 13 from the right by $(\mathcal{K}+\mathcal{J}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})$, we have:
$\mathbf{I}$ $=(\mathcal{Z}+\mathcal{J})(\mathcal{K}+\mathcal{J}-\tilde{\mathbf{D}}_{c}^{-1}+\mathbf{I})$ (14)
$=\mathcal{Z}\mathcal{K}+\mathcal{Z}\mathcal{J}-\mathcal{Z}\widetilde{\mathbf{D}}^{-1}+\mathcal{Z}+\mathcal{J}\mathcal{K}+\mathcal{J}^{2}-\mathcal{J}\widetilde{\mathbf{D}}^{-1}+\mathcal{J}.$
Since $\mathcal{J}^{2}=\mathcal{J}$, $\mathcal{J}\mathcal{K}=\mathcal{J}(\widetilde{\mathbf{D}}^{-1}-\mathbf{I})$, and $\mathcal{Z}\mathcal{J}=\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}\mathbf{Z}\mathbf{J}\widetilde{\mathbf{\Pi}}^{\frac{1}{2}}=\bm{0}_{n\times n}$, Eq. 14 can be simplified to:
$\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})=\mathbf{I}-\mathcal{J}.$ (15)
Similarly, by multiplying from the left, we establish that $(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}=\mathbf{I}-\mathcal{J}$. As $\pi^{\top}\mathbf{Z}=\bm{0}$, we have $\mathcal{J}\mathcal{Z}=0$. Thus, $\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}=\mathcal{Z}$. Furthermore, it can be straightforwardly inferred that $(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{J}=0$, then have $(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})=(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})$. Besides, considering the symmetry of the left part of Eq. 15, we have $(\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I}))^{\top}=\mathcal{Z}(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})$. Similarly, $((\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z})^{\top}=(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})\mathcal{Z}$. These derivations satisfy the sufficient conditions for the Moore–Penrose pseudoinverse, such that
$\mathcal{Z}=(\mathcal{K}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I})^{\dagger}.$ (16)
Finally, recovering $\mathcal{Z}$ and $\mathcal{K}$, which concludes the proof. ∎
A.3 SVD for $\mathcal{R}^{\dagger}$
Given a matrix $\mathcal{R}\in\mathbb{R}^{N\times N}$, its Moore-Penrose pseudoinverse can be directly computed with an SVD-based method. Specifically, we first perform truncated SVD on $\mathcal{R}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\widetilde{\mathcal{T}}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I}\approx\mathbf{U}_{q}\Sigma_{q}\mathbf{V}_{q}^{\top}$, where $\mathbf{U}_{q}\in\mathbb{R}^{N\times q}$ and $\mathbf{V}_{q}\in\mathbb{R}^{N\times q}$ contains the first $q$ columns of $\mathbf{U}$ and $\mathbf{V}$. $\Sigma_{q}\in\mathbb{R}^{q\times q}$ is the diagonal matrix of $q$ largest singular values. It is a $q$-rank approximation of $\mathcal{R}$, which holds that $\mathrm{rank}(\mathrm{R})=q$. Then the Moore-Penrose pseudoinverse of $\mathcal{R}$ can be easily computed as follows:
$\mathcal{R}^{\dagger}=\mathbf{U}_{q}\Sigma_{q}^{-1}\mathbf{V}_{q}^{\top}.$ (17)
To leverage sparsity of $\mathcal{R}$ to avoid $\mathcal{O}(N^{3})$ complexity, we adopt the randomized SVD algorithm proposed by (Halko et al., 2011; Cai et al., 2023) to first approximate the range of the input matrix with a low-rank orthonormal matrix, and then perform SVD on this smaller matrix:
$\hat{\bm{U}}_{q},\hat{\Sigma}_{q},\hat{\bm{V}}_{q}^{\top}=\operatorname{ApproxSVD}(\mathcal{R},q),\quad\hat{\mathcal{R}}_{SVD}=\hat{\bm{U}}_{q}\hat{\Sigma}_{q}\hat{\bm{V}}_{q}^{\top},$ (18)
where $\hat{\bm{U}}_{q}$, $\hat{\Sigma}_{q}$, and $\hat{\bm{V}}_{q}$ are the approximated versions of $\mathbf{U}_{q}$, $\Sigma_{q}$, and $\mathbf{V}_{q}$. Then the Moore-Penrose pseudoinverse of $\mathcal{R}$ can be computed by:
$\mathcal{R}^{\dagger}=\hat{\bm{U}}_{q}\hat{\Sigma}^{-1}_{q}\hat{\bm{V}}_{q}^{\top}.$ (19)
The computation cost of randomized truncated SVD takes $\mathcal{O}(qK)$, where $K$ is the number of non-zero elements in $\mathcal{R}$, so we have $K=|E|$. Thus, the sparsity degree of $\mathcal{R}$ can determine the time complexity of its Moore-Penrose pseudoinverse, which demonstrates the importance of Lemma 5.1.